OpenStudy (superhelp101):

When 190 grams of titanium tetrachloride react with an excess of water, as shown in the unbalanced chemical equation below, how many moles of hydrochloric acid will be produced? TiCl4 + H2O yields TiO2 + HCl 0.25 moles 0.5 moles 2.0 moles 4.0 moles

OpenStudy (superhelp101):

@asib1214

OpenStudy (asib1214):

balance the equation first.

OpenStudy (superhelp101):

@Abmon98 is this like the last question???

OpenStudy (abmon98):

Yes its the same method why dont you try solving it

OpenStudy (superhelp101):

balence equation2N2H4 + N2O4 + => 3N2 + 4H2O

OpenStudy (abmon98):

TiCl4 + H2O yields TiO2 + HCl On the Reactant Side: Ti:1 Cl:4 H:2 O:1 On the Product Side: Ti:1 Cl:1 H:1 O:2 Add a coefficient of 4 to HCl TiCl4 + H2O yields TiO2 + 4HCl On the Reactant Side: Ti:1 Cl:4 H:2 O:1 On the Product Side: Ti:1 Cl:4 H:4 O:2 Add a Coefficient of 2 to H2O TiCl4 + 2H2O yields TiO2 + 4HCl On the Reactant Side: Ti:1 Cl:4 H:4 O:2 On the Product Side: Ti:1 Cl:4 H:4 O:2 Number of Moles=Mass(g)/Molar Mass(g/mol) 190/189.679=1.00 mole Look at the Mole Ratio of TiCl4:HCL 1:4 1:x moles of HCL x-1*4/1=4moles

OpenStudy (abmon98):

2N2H4 + N2O4 + => 3N2 + 4H2O On The Reactant Side: N:6 H:8 O:4 On The Product Side: N :6 H:8 O:4 So this chemical equation is balanced

OpenStudy (superhelp101):

cool! Thanks for explaining it again I am glad i got the balance equation right though!

OpenStudy (abmon98):

Your welcome :)